# Resolving Multi-Condition Confusion for Finetuning-Free Personalized Image Generation

### v1

@@ Abstract
Personalized text-to-image generation methods can generate customized images based on the reference images, which have garnered wide research interest. Recent methods propose a finetuning-free approach with a decoupled cross-attention mechanism to generate personalized images requiring no test-time finetuning. However, when multiple reference images are provided, the current decoupled cross-attention mechanism encounters the object confusion problem and fails to map each reference image to its corresponding object, thereby seriously limiting its scope of application. To address the object confusion problem, in this work we investigate the relevance of different positions of the latent image features to the target object in diffusion model, and accordingly propose a weighted-merge method to merge multiple reference image features into the corresponding objects. Next, we integrate this weighted-merge method into existing pre-trained models and continue to train the model on a multi-object dataset constructed from the open-sourced SA-1B dataset. To mitigate object confusion and reduce training costs, we propose an object quality score to estimate the image quality for the selection of high-quality training samples. Furthermore, our weighted-merge training framework can be employed on single-object generation when a single object has multiple reference images. The experiments verify that our method achieves superior performance to the state-of-the-arts on the Concept101 dataset and DreamBooth dataset of multi-object personalized image generation, and remarkably improves the performance on single-object personalized image generation. Our code is available at  https://github.com/hqhQAQ/MIP-Adapter.
@@ 1 Introduction
Early finetuning-free methods, such as InstantBooth [17] and FastComposer [19], simply integrate the features of the reference image into the text embeddings and feed them into the text encoder, without fully exploiting the information from the reference image. Recent finetuning-free methods, such as IP-Adapter [21], more comprehensively utilize the features of the reference image by training additional cross-attention layers to integrate reference image features into the intermediate layers of the diffusion model, and achieve comparable performance to the finetuning-based methods. However, the current decoupled cross-attention only considers one reference image for each generation. When multiple reference images are provided, the decoupled cross-attention suffers from the object confusion problem if applied straightforwardly, wherein object features in the reference images are assigned to the wrong objects in the generated images, as illustrated in Figure 1 (A). Some previous image generation methods [20] attempt to mitigate the object confusion issue by incorporating the object features into the corresponding regions of latent image features in the diffusion model. Nevertheless, as the object information is distributed over the entire image feature space rather than confined to the corresponding local region owing to large receptive fields in deep networks [12, 1], the generated images can be limited in faithfulness to the reference images (i.e., the appearance differs between the generated and the reference images), as shown in Figure 2.
@@ 4 Experiments
Implementation details. Our main experiments are conducted on the pre-trained IP-Adapter with sdxl model [14] and sdxl_plus model as the text-to-image diffusion models and OpenCLIP ViT-bigG/14 as the image encoder. The parameters of sdxl model and image encoder are frozen, and only the parameters for projecting image features and predicting text weights are trainable. During training, we adopt AdamW optimizer with a learning rate of 1e-4, and train the model on 8 PPUs for 30,000 steps with a batch size of 4 per PPU. To enable classifier-free guidance, we use a probability of 0.05 to drop text and image individually, and a probability of 0.05 to drop text and image simultaneously. During inference, we adopt DDIM sampler with 50 steps and set the guidance scale to 7.5. We also conduct experiments on other pre-trained models based on decoupled cross-attention to verify the generalization ability of our method, in S2.2 of the appendix.

### v2

@@ Abstract
Personalized text-to-image generation methods can generate customized images based on the reference images, which have garnered wide research interest. Recent methods propose a finetuning-free approach with a decoupled cross-attention mechanism to generate personalized images requiring no test-time finetuning. However, when multiple reference images are provided, the current decoupled cross-attention mechanism encounters the object confusion problem and fails to map each reference image to its corresponding object, thereby seriously limiting its scope of application. To address the object confusion problem, in this work we investigate the relevance of different positions of the latent image features to the target object in diffusion model, and accordingly propose a weighted-merge method to merge multiple reference image features into the corresponding objects. Next, we integrate this weighted-merge method into existing pre-trained models and continue to train the model on a multi-object dataset constructed from the open-sourced SA-1B dataset. To mitigate object confusion and reduce training costs, we propose an object quality score to estimate the image quality for the selection of high-quality training samples. Furthermore, our weighted-merge training framework can be employed on single-object generation when a single object has multiple reference images. The experiments verify that our method achieves superior performance to the state-of-the-arts on multi-object personalized image generation, and remarkably improves the performance on single-object personalized image generation. Our code is available at  https://github.com/hqhQAQ/MIP-Adapter.
11footnotetext: * Equal contribution.22footnotetext: $\dagger$ Corresponding author.
@@ 1 Introduction
Early finetuning-free methods, such as InstantBooth [18] and FastComposer [20], simply integrate the features of the reference image into the text embeddings and feed them into the text encoder, without fully exploiting the information from the reference image. Recent finetuning-free methods, such as IP-Adapter [22], more comprehensively utilize the features of the reference image by training additional cross-attention layers to integrate reference image features into the intermediate layers of the diffusion model, and achieve comparable performance to the finetuning-based methods. However, the current decoupled cross-attention only considers one reference image for each generation. When multiple reference images are provided, the decoupled cross-attention suffers from the object confusion problem if applied straightforwardly, wherein object features in the reference images are assigned to the wrong objects in the generated images, as illustrated in Figure 1. Some previous image generation methods [21] attempt to mitigate the object confusion issue by incorporating the object features into the corresponding regions of latent image features in the diffusion model. Nevertheless, as the object information is distributed over the entire image feature space rather than confined to the corresponding local region owing to large receptive fields in deep networks [13, 1], the generated images can be limited in faithfulness to the reference images (i.e., the appearance differs between the generated and the reference images), as shown in Figure 2.
@@ 4 Experiments
Implementation details. Our main experiments are conducted on the pre-trained IP-Adapter with sdxl model [15] and sdxl_plus model [5] as the text-to-image diffusion models and OpenCLIP ViT-bigG/14 as the image encoder. The parameters of sdxl & sdxl_plus model and image encoder are frozen, and only the parameters for projecting image features and predicting text weights are trainable. During training, we adopt AdamW optimizer with a learning rate of 1e-4, and train the model on 8 PPUs for 30,000 steps with a batch size of 4 per PPU. To enable classifier-free guidance, we use a probability of 0.05 to drop text and image individually, and a probability of 0.05 to drop text and image simultaneously. During inference, we adopt DDIM sampler with 50 steps and set the guidance scale to 7.5. We also conduct experiments on other pre-trained models based on decoupled cross-attention to verify the generalization ability of our method, in S2.2 of the appendix.